# Leveraging MLLM Embeddings and Attribute Smoothing for Compositional Zero-Shot Learning

### v1

@@ Abstract
Compositional zero-shot learning (CZSL) aims to recognize novel compositions of attributes and objects learned from seen compositions. Previous works disentangle attribute and object by extracting shared and exclusive parts between image pairs sharing the same attribute (object), as well as aligning them with pretrained word embeddings to improve unseen attribute-object recognition. Despite the significant achievements of existing efforts, they are hampered by three limitations: (1) the efficacy of disentanglement is compromised due to the influence of the background and the intricate entanglement of attribute with object in the same parts. (2) existing word embeddings fail to capture complex multimodal semantic information. (3) overconfidence exhibited by existing models in seen compositions hinders their generalization to novel compositions. Being aware of these, we propose a novel framework named Multimodal Large Language Model (MLLM) embeddings and atTRibute smoothIng guiDEd diseNTanglement (TRIDENT) for CZSL. First, we leverage feature adaptive aggregation modules to mitigate the impact of background, and utilize learnable condition masks to capture multigranularity features for disentanglement. Then, the last hidden states of MLLM are employed as word embeddings for their superior representation capabilities. Moreover, we propose attribute smoothing with auxiliary attributes generated by Large Language Model (LLM) for seen compositions, addressing the issue of overconfidence by encouraging the model to learn more attributes in one given composition. Extensive experiments demonstrate that TRIDENT achieves state-of-the-art performance on three benchmarks.
@@ 1 Introduction
As for the study of compositional generalization ability inherent to humans, compositional zero-shot learning (CZSL) [22, 26, 33] is proposed to enable machines to recognize unseen attribute-object compositions by leveraging knowledge of attributes and objects (i.e., primitives) learned from seen compositions. Specifically, in the training phase, models are provided with images and compositional labels (e.g., ripe orange and peeled apple). During the testing phase, given an image depicting a novel composition (e.g., peeled orange), models are assigned to classify the image into the corresponding category [44].
Prior works [22, 27] focus on mapping the visual features and the word embeddings of compositions into a joint space. These methods have poor generalization capability to unseen compositions, as they fail to learn primitives. Therefore, recent studies [38, 8, 14] consider visual disentanglement. Among them, some prominent works deploy a triplet of images to disentangle: a given image (noted as the main image), and two supplementary images, each sharing either the same attribute or the same object as the main image. The triplet of images is treated as two image pairs for subsequent analysis. These approaches aim to disentangle attribute and object by analyzing the shared and exclusive features of the image pair, as well as aligning them with word embeddings (e.g., GloVe [32]), as shown in Figure 1. Although these pioneer research studies have achieved great progress, they exhibit three limitations:
L1: Disentanglement is impeded due to the influence of the background and the intricate entanglement of attribute with object in the same parts of image. On the one hand, models tend to extract the background feature unique to one image in the pair as the disentangled exclusive features. On the other hand, some existing methods [36, 38] compute the similarity of image pairs for disentanglement at the spatial level. However, disentangling attribute and object at the spatial level presents significant challenges because they entangle in the same spatial features. Taking an image of ripe apple as an example, the spatial regions corresponding to the attribute ”ripe” and the object ”apple” are fully co-located.
@@
L2: Existing word embeddings lack the depth needed to capture complex multimodal semantic information. To begin with, word embeddings, such as Word2Vec [21] and GloVe [32], are grounded in word frequency and contextual co-occurrence, rather than capturing high-level semantic nuances[39]. Moreover, the process of aligning visual features with word embeddings can be viewed as a form of cross-modal matching; however, these word embeddings are trained only in a single text modal, failing to capture cross-modal information between images and texts.
L3: Existing methods display excessive confidence in seen compositions, impairing their ability to generalize toward novel compositions. Due to the one-hot label used during training, these approaches are limited by learning only one attribute and object, neglecting the fact that objects naturally exhibit multiple attributes [43]. Consequently, models exhibit overconfidence in the disentangled ground-truth attribute, treating other attributes that can describe the object as negative attributes, which results in the diminished performance on unseen compositions.
To address the aforementioned limitations, we propose a novel framework named Multimodal Large Language Model (MLLM) embeddings and atTRibute smoothIng guiDEd diseNTanglement (TRIDENT), which consists of three major modules: visual feature extraction, attribute-object disentanglement, and feature alignment. The first module leverages feature adaptive aggregation (FAA) modules to mitigate the impact of background noise, and exploits learnable condition masks to learn multi-granularity features to improve subsequent disentanglement. The second module aims at leveraging shared and exclusive weights of image pairs to disentangle attribute and object under the the paradigm that apart from the shared features of the image pair, each image has its own exclusive features. The third module is intended to align the visual features of compositions and disentangled primitives with the last hidden states of an MLLM, LLaVA v1.5 [16], i.e., MLLM embeddings. This is inspired by some works [41, 23, 24, 12], which find that the last hidden states of (M)LLM exhibit powerful representational capabilities in embedding tasks, such as retrieval and classification. Moreover, to tackle the issue that the ineffective overconfidence of the models regarding ground-truth attribute hinders them from generalizing to unseen compositions, we exploit a Large Language Model (LLM), GPT-3.5 [29] to generate auxiliary attributes based on attribute-object compositions and perform label smoothing for attributes, i.e., attribute smoothing.
@@
1. We propose novel feature adaptive aggregation modules to reduce the impact of background, and utilize learnable condition masks to capture multi-granularity features for disentanglement in CZSL.
2. We employ both LLM and MLLM to guide attribute-object disentanglement by generating auxiliary attributes and representing word embeddings, respectively. To the best of our knowledge, we are the first to leverage both LLM and MLLM to advance disentanglement in CZSL task.
3. We conduct extensive experiments to evaluate our method on three CZSL benchmarks, showing that TRIDENT has achieved state-of-the-art performance. The source code will be released soon  [^1] https://github.com/xud-yan/Trident.
@@ 2 Related Work
Compositional zero-shot learning (CZSL). Prior works in CZSL can be broadly divided into two main streams. One main stream is to learn representations of compositions in a joint space. SymNet [15] proposes to learn symmetry property in compositions. Co-CGE [20] leverages a Graph Convolutional Neural Network to learn compositional representations. The other main stream aims at disentangling visual representations of primitives to reduce composition learning into primitive learning. SCEN [13] leverages contrastive loss to excavate discriminative prototypes of primitives. OADis [38] disentangles primitives by affinity modules. CANet [42] learns conditional attribute conditioned on the recognized object and the input image.
More recent works [28, 18, 9] focus on leveraging the encyclopedic knowledge of pretrained vision-language models (VLM), such as Contrastive Language-Image Pre-training (CLIP) [35] and Context Optimization (CoOp) [46], to encode and align images and texts.
Large language model (LLM). LLMs have realized significant advancements thanks to the scaling up of training data and the increase in the number of parameters. Early models, such as BERT [6] and GPT-2 [34], initially exhibit strong capabilities in understanding and generating human-like language. Subsequently, GPT-3 [4] and LLaMA [40] demonstrate great breakthroughs across numerous language benchmarks. Moreover, by performing instruction fine-tuning on LLM, ChatGPT [30] and Vicuna [5, 45] are able to comprehend and follow human instructions better.
Expanding on LLM, Multimodal Large Language Model (MLLM) incorporates a pretrained visual encoder for vision-language tasks. Flamingo [1] first integrates Vision Transformer (ViT) [7] and LLM by gated cross-attention. Recently, LLaVA [17] and LLaVA v1.5 [16] introduce visual instruction tuning to enhance instruction following capability. The visual understanding part of LLaVA v1.5 consists of a ViT and a multilayer perceptron (MLP) cross-modal connector (CMC). CMC processes visual features before the last layer of ViT, aligning the visual space of ViT with the linguistic space of LLM. We choose LLaVA v1.5 as our foundational MLLM as it has demonstrated state-of-the-art performance across various tasks.
Recently, exploring the powerful language capabilities of (M)LLM to handle representation tasks (e.g., retrieval) has emerged as a prominent research domain. SGPT [23] exploits the last hidden states of LLM for the input token sequence or a special learnable token to derive representational embeddings. Subsequently, GritLM [24] applies mean pooling over the last hidden states of LLM to yield the textual embeddings. FROMAGe [12] uses a learnable token to represent the text fed into MLLM for image retrieval.
Figure 2: The overall architecture of our proposed TRIDENT. TRIDENT consists of three major modules: (a) visual feature extraction, (b) attribute-object disentanglement, and (c) feature alignment.
@@ 3.1 Task Formulation
Compositional zero-shot learning (CZSL) aims at learning a model that can recognize unseen compositions of attributes and objects that are learned from seen compositions. Given an attribute set $\mathcal{A}$ and an object set $\mathcal{O}$, the attributes and objects are composed to form a composition set $\mathcal{C}=\mathcal{A}\times\mathcal{O}$. The composition set $\mathcal{C}$ is divided into two disjoint sets: the seen composition set $\mathcal{C}_{s}$ and the unseen composition set $\mathcal{C}_{u}$, where $\mathcal{C}_{s}\cap\mathcal{C}_{u}=\varnothing$ and $\mathcal{C}_{s}\cup\mathcal{C}_{u}=\mathcal{C}$. The model is trained with a seen training set $\mathcal{D}_{tr}=\{(x_{s},c_{s})\}$, where $x_{s}\in\mathcal{X}_{s}$ is an image from the seen image set $\mathcal{X}_{s}$ corresponding to the seen composition set $\mathcal{C}_{s}$, and $c_{s}\in\mathcal{C}_{s}$ is the label of $x_{s}$. Following the Generalized CZSL [33], the model is evaluated on a predefined test set $\mathcal{D}_{te}=\{(x_{te},c_{te})\}$, where $x_{te}\in\mathcal{X}_{te}$ is an image from the unseen image set $\mathcal{X}_{te}$ corresponding to the composition subset $\mathcal{C}_{te}$ of $\mathcal{C}$, i.e., $\mathcal{C}_{te}\subseteq\mathcal{C}$, and $c_{te}\in\mathcal{C}_{te}$ is the label of $x_{te}$. The aim of CZSL task is to learn a model $M:\mathcal{X}_{te}\rightarrow\mathcal{C}_{te}$ that predicts labels $c_{te}$ from $\mathcal{C}_{te}$ for the input images $x_{te}\in\mathcal{X}_{te}$.
@@ 3.2 TRIDENT
As the major novelty, we propose a novel framework named MLLM embeddings and attribute smoothing guided disentanglement framework (TRIDENT) for CZSL, as shown in Figure 2. It consists of three major modules: (1) visual feature extraction, (2) attribute-object disentanglement, and (3) feature alignment. We now detail each module of TRIDENT in this section.
@@ 3.2.1 Visual Feature Extraction
As shown in Figure 2, we denote a given image with the attribute-object composition label (e.g. ripe apple) as the main image $x^{m}$, and randomly sample an image with the same attribute $x^{a}$ (i.e., ripe orange), as well as an image sharing the same object $x^{o}$ (i.e., peeled apple) to comprise a triplet image set. For the convenience of expression, we simply use $x^{img}$ (where $img\in\{m,a,o\}$) to collectively denote the images as they are processed using the same module.
Visual feature extraction backbone. As mentioned before, since LLaVA v1.5 is used as our fundamental MLLM, we directly leverage the visual encoder, ViT, and cross-modal connector (CMC) from the model to extract visual features. Specifically, the image $x^{img}$ is partitioned into $n$ patch tokens, which are subsequently put into ViT along with the [CLS] token. Afterward, the output of patch tokens before the last layer of ViT is fed into the CMC module, as implemented in LLaVA v1.5. To align the dimension of patch tokens output by CMC with that of [CLS] token produced by ViT, the patch tokens output by CMC are input into a linear layer. Consequently, we obtain one feature vector of [CLS] token $f^{img}_{cls}\in\mathbb{R}^{d}$ and a patch feature matrix of $n$ patch tokens $F^{img}_{patch}\in\mathbb{R}^{n\times d}$, where $d$ is the dimension of the features.
Local features extraction. Intuitively, the composition (e.g., ripe apple) only occupies a few parts of the image. Since each patch token usually corresponds to one local region of the image, to filter out background noise and focus on related regions, we deploy a set of feature adaptive aggregation (FAA) modules to derive $p$ relevant local features of $x^{img}$, where each FAA module is formulated as follows:
$\left\{\begin{array}[]{ll}v=agg\otimes F^{img}_{patch}\\[5.0pt] agg=\sigma(Conv(F^{img}_{patch}))\end{array}\right.$ (1)
where $Conv(\cdot)$ represents the 1 $\times$ 1 convolution layer, $\sigma(\cdot)$ denotes the sigmoid activation function, $agg\in\mathbb{R}^{n}$ is the weight vector, the $k$-th element of $agg$ is the weight for $k$-th patch feature. $\otimes$ represents matrix product, and $v\in\mathbb{R}^{d}$ is the local feature obtained by an FAA module. We vertically concatenate the local features produced by $p$ FAA modules to obtain the local feature matrix $F^{img}_{l}\in\mathbb{R}^{p\times d}$.
Global features extraction. Normally, the ViT output of [CLS] token is regarded as containing various global information of the image, which highly entangles both attribute and object features together[8]. To disperse multi-granularity global information into different representations, $q$ learnable condition masks are applied to $f^{img}_{cls}$ to obtain $q$ different global representations, where each global representation is computed as:
@@
where $u\in\mathbb{R}^{d}$ denotes each global representation. Here $c\in\mathbb{R}^{d}$ refers to each learnable condition mask and $\odot$ is the element-wise multiplication. Consequently, we vertically concatenate $q$ global representations to derive the global feature matrix $F^{img}_{g}\in\mathbb{R}^{q\times d}$.
Features concatenation. Finally, $F^{img}_{l}$ and $F^{img}_{g}$ are vertically concatenated to form the visual features of $x^{img}$, i.e., $F^{img}=[F^{img}_{l},F^{img}_{g}]\in\mathbb{R}^{(p+q)\times d}$, which is used for the following disentanglement of attribute and object.
Orthogonal regularization. We ideally want features extracted by different modules can represent different information of the image $x^{img}$. To this end, we further introduce the orthogonal regularization, i.e.:
$\mathcal{L}_{ortho}=\sum_{img\in\{m,a,o\}}(||F^{img}{F^{img}}^{T}-I||_{Fro})$ (3)
where $I\in\mathbb{R}^{(p+q)\times(p+q)}$ is the identity matrix. $||\cdot||_{Fro}$ refers to the Frobenius norm of the matrix.
Image embedder. Inspired by [26], for the input image $x^{img}$, we first use AveragePools $Avg(\cdot)$ on $F^{img}_{g}$ and $F^{img}_{l}$, respectively, and horizontally concatenate them by $Cat(\cdot,\cdot)$ to aggregate both global and local visual information of $x^{img}$ corresponding to the composition label. Then the concatenated feature passes through a linear layer $Lin_{comp}(\cdot)$ to derive the final feature representation $f^{img}_{comp}$. This module is formulated as follows:
@@
where $f^{img}_{comp}\in\mathbb{R}^{2d}$ denotes the visual feature corresponding to the composition. This module is designed to learn the visual features of images associated with their corresponding composition labels, serving as the primary branch for inference.

### v2

@@ Abstract
Compositional zero-shot learning (CZSL) aims to recognize novel compositions of attributes and objects learned from seen compositions. Previous works disentangle attributes and objects by extracting shared and exclusive parts between the image pair sharing the same attribute (object), as well as aligning them with pretrained word embeddings to improve unseen attribute-object recognition. Despite the significant achievements of existing efforts, they are hampered by three limitations: (1) The efficacy of disentanglement is compromised due to the influence of the background and the intricate entanglement of attributes with objects in the same parts. (2) Existing word embeddings fail to capture complex multimodal semantic information. (3) Overconfidence exhibited by existing models in seen compositions hinders their generalization to novel compositions. Being aware of these, we propose a novel framework named multimodal large language model (MLLM) embeddings and attribute smoothing guided disentanglement for CZSL. First, we leverage feature adaptive aggregation modules to mitigate the impact of background, and utilize learnable condition masks to capture multi-granularity features for disentanglement. Moreover, the last hidden states of MLLM are employed as word embeddings for their superior representation capabilities. Furthermore, we propose attribute smoothing with auxiliary attributes generated by the large language model (LLM) for seen compositions to address the overconfidence challenge. Extensive experiments demonstrate that our method achieves state-of-the-art performance on three challenging datasets. The supplementary material and source code will be available at https://github.com/xud-yan/Trident.
@@ 1 Introduction
As for the study of compositional generalization ability inherent in human beings, compositional zero-shot learning (CZSL) Misra et al. (2017); Nagarajan and Grauman (2018); Purushwalkam et al. (2019) is proposed to enable machines to recognize unseen compositions by leveraging knowledge of attributes and objects (i.e., primitives) learned from seen compositions. Specifically, in the training phase, the models are provided with images of seen compositions (e.g., ripe orange and peeled apple). During the testing phase, given an image that depicts a novel composition (e.g., peeled orange), models are assigned to correctly recognize it Zhang et al. (2022).
Prior works Misra et al. (2017); Nan et al. (2019); Naeem et al. (2021) focus on mapping the visual features and the word embeddings of compositions into a joint space. These methods have poor generalization capabilities on unseen compositions due to the recombination of primitives. Therefore, recent studies Saini et al. (2022); Wang et al. (2023); Li et al. (2024); Zhang et al. (2024) consider visual disentanglement. Among them, some prominent works Hao et al. (2023) deploy a triplet of images to disentangle visual features: a given image and two supplementary images, each sharing either the same attribute or object as the given image. The triplet of images is treated as two image pairs for subsequent analysis. These approaches aim to disentangle attribute and object by extracting the shared and exclusive features of the image pair, as well as aligning them with word embeddings (e.g., GloVe Pennington et al. (2014)), as shown in Figure 1. Although these pioneering research studies have made great progress, they exhibit three limitations:
@@
L1: Disentanglement is impeded due to the influence of the background and the intricate entanglement between attributes and objects in the same parts of images. On the one hand, models tend to extract the background features unique to one image in the pair as the disentangled exclusive features. On the other hand, some existing methods Ruis et al. (2021); Saini et al. (2022) compute the similarity of the image pair for disentanglement at the spatial level. However, this paradigm is limited by the frequent entanglement of attributes and objects within the same image regions.
L2: Existing word embeddings lack the depth needed to capture complex multimodal semantic information. To begin with, word embeddings (e.g., GloVe Pennington et al. (2014)) are grounded in word frequency and contextual co-occurrence, overlooking the high-level semantic nuances Sarzynska-Wawer et al. (2021). Moreover, the process of aligning visual features with word embeddings can be viewed as a form of cross-modal matching; however, these word embeddings are trained only in a single text modality, failing to capture multimodal information between images and texts.
L3: Existing methods display excessive confidence in seen compositions, impairing their ability to generalize towards novel compositions. Specifically, due to the one-hot label used during training, these approaches are limited by learning only one disentangled attribute and object, neglecting the fact that objects naturally exhibit multiple attributes Xu et al. (2024). Consequently, models exhibit overconfidence in the disentangled ground-truth attribute, treating other attributes that can describe the object as negative ones, which results in the diminished generalization to unseen compositions.
Being aware of these, we propose a novel framework named multimodal large language model (MLLM) embeddings and atTRibute smoothIng guiDEd diseNTanglement (TRIDENT), which consists of three major modules: visual feature extraction, attribute-object disentanglement, and feature alignment. The first module leverages feature adaptive aggregation (FAA) modules to mitigate the impact of background noise, and exploits learnable condition masks for multi-granularity feature learning at the dimensional level to improve subsequent disentanglement. The second module aims at leveraging shared and exclusive weights of image pairs to disentangle attributes and objects under the paradigm that apart from the shared features of the image pair, each image has its own exclusive features. The third module is intended to align the visual features of compositions and disentangled primitives with the last hidden states of MLLM (i.e., MLLM embeddings). This is inspired by prior works Wang and Kuo (2020); Muennighoff (2022); Muennighoff et al. (2024), which reveal that the last hidden states of (M)LLM exhibit powerful representation capabilities in embedding tasks (e.g., retrieval and classification). Moreover, to tackle the issue that the overconfidence of the models regarding the ground-truth attribute hinders them from generalizing to unseen compositions, we exploit the large language model (LLM) to generate auxiliary attributes for compositions and perform label smoothing for attributes (i.e., attribute smoothing).
@@
1. We propose novel feature adaptive aggregation modules to reduce the impact of background, and utilize learnable condition masks to capture multi-granularity features at the dimensional level for disentanglement in CZSL.
2. We employ both LLM and MLLM to guide attribute-object disentanglement by generating auxiliary attributes and representing primitive words for CZSL, respectively.
3. Extensive experiments conducted on three challenging datasets (MIT-States Isola et al. (2015), C-GQA Naeem et al. (2021), and VAW-CZSL Saini et al. (2022)) show that TRIDENT has achieved state-of-the-art performance.
@@ 2 Related Work
Figure 2: The overall architecture of our proposed TRIDENT. The model consists of three major modules: (a) visual feature extraction, (b) attribute-object disentanglement, and (c) feature alignment.
Compositional zero-shot learning (CZSL). Prior works in CZSL can be broadly divided into two main streams. One main stream is learning visual representations and textual labels of compositions in a joint space. SymNet Li et al. (2020) aims to learn the symmetry property in compositions. Co-CGE Mancini et al. (2022) leverages a graph convolutional neural network to learn composition representations. The other main stream aims at disentangling visual representations of primitives to reduce composition learning to primitive learning. SCEN Li et al. (2022) leverages contrastive loss to excavate discriminative prototypes of primitives. CANet Wang et al. (2023) learns the conditional attribute conditioned on the recognized object and the input image. More recent works Nayak et al. (2023); Lu et al. (2023); Huang et al. (2024) leverage the encyclopedic knowledge of pre-trained vision-language models (VLM) like CLIP Radford et al. (2021) to encode and align images and texts.
Large language model (LLM). LLMs have realized significant advancements thanks to the scaling up of training data and the increase in the number of parameters. Early models, such as GPT-2 Radford et al. (2019), initially exhibit strong capabilities in understanding and generating human-like language. Subsequently, GPT-3 Brown et al. (2020) and LLaMA Touvron et al. (2023) demonstrate great breakthroughs across numerous language benchmarks.
Expanding on LLMs, multimodal large language models (MLLM) incorporate a visual encoder for vision-language tasks. Flamingo Alayrac et al. (2022) integrates Vision Transformer (ViT) Dosovitskiy et al. (2021) and LLM by gated cross-attention. LLaVA Liu et al. (2024b) and LLaVA v1.5 Liu et al. (2024a) introduce visual instruction tuning to enhance the instruction following capability. The visual understanding part of LLaVA v1.5 consists of a ViT and an MLP cross-modal connector. We choose LLaVA v1.5 as our foundational MLLM for its state-of-the-art performance.
Recently, exploring the powerful embedding capabilities of (M)LLM to handle representation tasks (e.g., retrieval) has emerged as a prominent research domain. SGPT Muennighoff (2022) exploits the last hidden states of LLM for the input token sequence or a special learnable token to derive representational embeddings. Subsequently, GritLM Muennighoff et al. (2024) applies mean pooling over the last hidden states of LLM to produce the textual embeddings.
@@ 3.1 Task Formulation
Compositional zero-shot learning (CZSL) aims at learning a model that can recognize unseen compositions of attributes and objects that are learned from seen compositions. Given an attribute set $\mathcal{A}$ and an object set $\mathcal{O}$, the attributes and objects are composed to form a composition set $\mathcal{C}=\mathcal{A}\times\mathcal{O}$. The composition set $\mathcal{C}$ is divided into two disjoint sets: the seen composition set $\mathcal{C}_{s}$ and the unseen composition set $\mathcal{C}_{u}$, i.e., $\mathcal{C}_{s}\cap\mathcal{C}_{u}=\varnothing$. The model learns from a seen training set $\mathcal{D}_{tr}=\{(x_{s},c_{s})\}$, where $x_{s}$ is an image of the seen composition label $c_{s}\in C_{s}$. Following the Generalized CZSL Purushwalkam et al. (2019), the model is evaluated on a predefined test set $\mathcal{D}_{te}=\{(x_{te},c_{te})\}$, where $x_{te}$ is a test image from the predefined composition subset $\mathcal{C}_{te}$ of $\mathcal{C}$, i.e., $\mathcal{C}_{te}\subseteq\mathcal{C}$, and $c_{te}\in\mathcal{C}_{te}$ is the label of $x_{te}$.
@@ 3.2 TRIDENT
As the major novelty, we propose a novel framework named MLLM embeddings and attribute smoothing guided disentanglement (TRIDENT), as shown in Figure 2. It consists of three modules: (1) visual feature extraction, (2) attribute-object disentanglement, and (3) feature alignment.
@@ 3.2.1 Visual Feature Extraction
As shown in Figure 2, we denote a given image with the attribute-object composition label (e.g. ripe apple) as the main image $x^{m}$, and randomly sample an image with the same attribute $x^{a}$ (i.e., ripe orange) and an image sharing the same object $x^{o}$ (i.e., peeled apple) to form a triplet image set. For convenience of expression, we simply use $x^{img}$ (where $img\in\{m,a,o\}$) to collectively denote the images as they are processed using the same module.
Visual backbone. As mentioned before, since LLaVA v1.5 is used as our fundamental MLLM, we directly leverage its visual encoder (i.e., ViT) and cross-modal connector (CMC) to extract visual features. Specifically, the image $x^{img}$ is partitioned into $n$ patch tokens, which are subsequently put into ViT along with the [CLS] token. Afterward, the output of patch tokens before the last layer of ViT is fed into the CMC module, as implemented in LLaVA v1.5. To align the dimensions of the patch tokens output by CMC and the [CLS] token produced by ViT, the patch tokens output by CMC is input into a linear layer. Consequently, we obtain one feature vector of [CLS] token $f^{img}_{cls}\in\mathbb{R}^{d}$ and a patch feature matrix of $n$ patch tokens $F^{img}_{patch}\in\mathbb{R}^{n\times d}$.
Local features extraction. Intuitively, the composition (e.g., ripe apple) only occupies a few parts of the image. Since each patch token corresponds to one local region of the image, to filter out background noise and focus on related regions, we deploy $p$ feature adaptive aggregation (FAA) modules to derive $p$ relevant local features of $x^{img}$, where each FAA module is formulated as follows:
$\left\{\begin{array}[]{ll}v=agg\otimes F^{img}_{patch}\\[5.0pt] agg=ReLU(Conv(F^{img}_{patch}))\end{array}\right.$ (1)
where $Conv(\cdot)$ represents the 1 $\times$ 1 convolution layer, $agg\in\mathbb{R}^{n}$ is the weight vector, and the $k$-th element of $agg$ is the weight for $k$-th patch feature. $\otimes$ represents matrix product, and $v\in\mathbb{R}^{d}$ is the local feature obtained by an FAA module. We vertically concatenate the local features produced by $p$ FAA modules to obtain the local feature matrix $F^{img}_{l}\in\mathbb{R}^{p\times d}$.
Global features extraction. Normally, the [CLS] token output by ViT is regarded as containing various global information of the image, which highly entangles both attribute and object together Hao et al. (2023). To disperse multi-granularity global information into different representations at the dimensional level, $q$ learnable condition masks are applied to $f^{img}_{cls}$ to obtain $q$ different global representations. Each global representation is computed as:
@@
where $u\in\mathbb{R}^{d}$ denotes the global representation, $c\in\mathbb{R}^{d}$ refers to the learnable condition mask and $\odot$ is the element-wise multiplication. We vertically concatenate $q$ global representations to derive the global feature matrix $F^{img}_{g}\in\mathbb{R}^{q\times d}$.
Features concatenation. Finally, $F^{img}_{l}$ and $F^{img}_{g}$ are vertically concatenated to form the visual features of $x^{img}$, i.e., $F^{img}=[F^{img}_{l},F^{img}_{g}]\in\mathbb{R}^{h\times d}$ (where $h=p+q$), which is used for the following attribute-object disentanglement.
Orthogonal regularization. Different features should capture distinct and complementary aspects of the image. Therefore, we introduce the orthogonal regularization, i.e.:
$\mathcal{L}_{ortho}=\frac{1}{|img|\cdot|i|}\sum_{\begin{subarray}{c}img\in\{m,a,o\},i\in\{g,l\}\end{subarray}}(||F^{img}_{i}{F^{img}_{i}}^{T}-I_{i}||_{Fro})$ (3)
where $I_{i}$ denotes the identity matrix, and $||\cdot||_{Fro}$ refers to the Frobenius norm of the matrix.
Image embedder. Inspired by Nagarajan and Grauman (2018), for the input image $x^{img}$, we first use the average pooling $Avg(\cdot)$ on $F^{img}_{g}$ and $F^{img}_{l}$, respectively, and horizontally concatenate them by $Cat(\cdot,\cdot)$ to aggregate both global and local features of $x^{img}$. Then the concatenated feature passes through a linear layer $Lin_{comp}(\cdot)$ to derive the final visual feature $f^{img}_{comp}$ that represents the composition. This module is formulated as follows: